What is maximum height in projectile motion.


1
Expert's answer
2021-02-12T16:17:19-0500

The maximum height of the projectile is the height at which the vertical velocity of the projectile becomes zero. The maximum height of the projectile can be found as follows:


"y_{max}=\dfrac{v_0^2sin^2\theta}{2g},"

here, "v_0" is the initial velocity of the projectile, "\theta" is the launch angle and "g" is the acceleration due to gravity.